OpenStudy (anonymous):

Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (a1234):

First, put that equation in slope-intercept form. 2x + 3y = 5 3y = -2x + 5 y = -2/3x + 5/3 Use y - y1 = m(x - x1) y + 3 = -2/3( x - 9) y + 3 = -2/3x + 6 y = -2/3x + 3

OpenStudy (anonymous):

Oh my gosh. Thank you so much :)

OpenStudy (a1234):

You're welcome.

OpenStudy (a1234):

Oh, I never knew that. The method I used above was the only one my textbook said. Thanks for informing me!